OpenStudy (anonymous):

Solve the following equation for d. Please show work. k=2c-9d/8

OpenStudy (anonymous):

k-2c=-9d/8 ...8(k-2c)=-9d ...(8(k-2c))/-9=d

OpenStudy (anonymous):

Take 9d/8 to the left side, and k over to the right to get...9d/8 = 2c-k...then multiply both by 8 to get 9d = 18c-8k, lastly divide by 9 to get d = (18c-8k)/9

OpenStudy (anonymous):

to get rid of the 8, multiply across by 8: 8k = 16c - 9d 8k - 16c = -9d -8/9k + 16c/9 = d

OpenStudy (anonymous):

d = (16c-8k)/9...hehe can't multiply